OpenStudy (anonymous):

How do you determine the inverse of this relation f(x)=1/(x-2)+3

jimthompson5910 (jim_thompson5910):

f(x)=1/(x-2)+3 y = 1/(x-2)+3 ... replace f(x) with y x = 1/(y-2)+3 ... swap x and y Now solve for y

jimthompson5910 (jim_thompson5910):

x = 1/(y-2)+3 x - 3 = 1/(y-2) (x - 3)(y-2) = 1 y-2 = 1/(x-3) y = 1/(x-3) + 2